# Peter Reinhard Hansen§    and   Chen Tong

### v3

@@ Appendix B Result for the Dynamic Models
B.1 MGF for ARG model
@@
×$\quad\times\exp\left(-\delta\log\left(1-\eta B_{1}(h,z)\right)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\sum_{j=1}^{p}\beta_{j}X_{T+1-j}\right)$
@@
B.2 The MGF in Heston-Nandi GARCH Model
We first have
×$\times\mathbb{E}_{t}\left[v_{2}\alpha z_{t+1}^{2}+\left(v_{1}-2v_{2}\alpha\gamma\right)\sqrt{h_{t+1}}z_{t+1}\right]$
Next, we show by induction that the conditional MGF has the form,
For $h=1$ we have
×$\quad\times\exp\left(\left(s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)}\right)h_{t+1}\right)$
$=\exp\left(A(h+1,s)+B(h+1,s)h_{t+1}\right),$
where
@@ B.2 The MGF in Heston-Nandi GARCH Model
An Direct Proof of Lemma A.1.
Proof of Lemma A.1. By the substitution, $y=zx\in\mathbb{C}$, we have
$\lim_{K\rightarrow\infty}\int_{0}^{K}x^{r}e^{-zx}\mathrm{d}x=\frac{1}{z^{r+1}}\int_{0}^{sK+itK}y^{r}e^{-y}\mathrm{d}y.$
The function $f(y)\equiv y^{r}e^{-y}$, has a pole at $y=0$ if $r<0$, and is a multi-valued function when $r$ is a non-integer. We will adopt the principal value by restricting the angle of $y=\lambda e^{i\theta}$ to $\theta\in(-\pi,\pi]$, and construct a closed contour integral that avoids the pole at $y=0$. The contour is illustrated in Figure S.1.
Figure S.1: The closed contour for the integral, $\oint y^{r}e^{-y}\mathrm{d}y$.
The function, $f(y)$, is analytic over the closed contour for $\lambda>0$, (i.e. it does not include any singularity). Therefore, by the Cauchy Integral Theorem, we have
$0=\oint f(y)\mathrm{d}y$ $=\int_{AB}f(y)\mathrm{d}y+\int_{BC}f(y)\mathrm{d}y+\int_{CD}f(y)\mathrm{d}y+\int_{DA}f(y)\mathrm{d}y,$
where the integral we seek is $-\int_{CD}f(y)\mathrm{d}y$ as $\lambda\rightarrow 0$ and $K\rightarrow\infty$.
For the integral over the arc, DA, we have $y=\lambda e^{i\theta}=\lambda\left(\cos\theta+i\sin\theta\right)$, and $dy=i\lambda e^{i\theta}d\theta$, such that
$\left|I_{DA}\right|=\left|\int_{DA}y^{r}e^{-y}\mathrm{d}y\right|$ $=\left|\int_{\pi/2}^{0}\lambda^{r}e^{i\theta r}e^{-\lambda\left(\cos\theta+i\sin\theta\right)}i\lambda e^{i\theta}\mathrm{d}\theta\right|$
$\leq\left|\lambda^{r+1}\right|e^{|\lambda|}\left|\int_{\pi/2}^{0}e^{i\theta r}e^{-i\lambda\sin\theta}ie^{i\theta}\mathrm{d}\theta\right|$
$\leq\left|\lambda^{r+1}\right|e^{|\lambda|}\left|\int_{\pi/2}^{0}\mathrm{d}\theta\right|=\tfrac{\pi}{2}\lambda^{{\rm Re}\left(r\right)+1}e^{|\lambda|}\rightarrow 0,\quad\text{as }\lambda\rightarrow 0\text{ for }{\rm Re}\left(r\right)>-1.$
For the BC-integral, we have $y=sK+iuK$ and $\mathrm{d}y=iK\mathrm{d}u$, such that
$\left|I_{BC}\right|=\left|\int_{sK}^{sK+itK}y^{r}e^{-y}\mathrm{d}y\right|$ $=\left|iK\int_{0}^{t}\left(sK+iuK\right)^{r}e^{-\left(sK+iuK\right)}\mathrm{d}u\right|$
$=\left|iK^{r+1}e^{-sK}\int_{0}^{t}\left(s+iu\right)^{r}e^{-iuK}\mathrm{d}u\right|$
$\leq K^{{\rm Re}\left(r\right)+1}e^{-sK}\left|\int_{0}^{t}\left|\left(s+iu\right)^{r}\right|\mathrm{d}u\right|\rightarrow 0,\qquad\text{as }K\rightarrow\infty,$
because $e^{-sK}$ decays faster than $K^{r+1}$ increases. So, as $K\rightarrow\infty$ and $\lambda\rightarrow 0$, we have
$0=\lim_{K\rightarrow\infty}\lim_{\lambda\rightarrow 0}\left(\int_{AB}f(y)\mathrm{d}y+\int_{CD}f(y)\mathrm{d}y\right)=\lim_{K\rightarrow\infty}\left(\int_{0}^{sK}f(y)\mathrm{d}y+\int_{sK+itK}^{0}f(y)\mathrm{d}y\right).$
The integral we seek is therefore given by,
$\lim_{K\rightarrow\infty}\int_{0}^{sK+itK}f(y)\mathrm{d}y=\lim_{K\rightarrow\infty}\int_{0}^{sK}f(y)\mathrm{d}y=\Gamma(r+1).$
This completes the proof. $\square$
@@
[…]

### v4

@@ Appendix B Result for the Dynamic Models
The MGF in Heston-Nandi GARCH Model
Proof of Proposition 1. We first have
$\times\mathbb{E}_{t}\left[v_{2}\alpha z_{t+1}^{2}+\left(v_{1}-2v_{2}\alpha\gamma\right)\sqrt{h_{t+1}}z_{t+1}\right]$
Next, we show by induction that the conditional MGF has the form,
For $h=1$ we have
$\quad\times\exp\left(\left(s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)}\right)h_{t+1}\right)$
$=\exp\left(A(h+1,s)+B(h+1,s)h_{t+1}\right),$
where
MGF for ARG model
@@
$\quad\times\exp\left(-\delta\log\left(1-\eta B_{1}(h,z)\right)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\sum_{j=1}^{p}\beta_{j}X_{T+1-j}\right)$
@@
Equation A.1.
A proof of Consider the case where $z$ is real and positive, i.e. $z=s>0$. From the definition of the gamma function and the substitution, $y\mapsto sx$, we have $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, which proves the identity for any $z\in\mathbb{R}^{+}$. Both sides of (A.1) are holomorphic functions for $\operatorname{Re}(z)>0$, and since they agree on $z\in\mathbb{R}^{+}$, it follows by the identity theorem that they agree for all $\operatorname{Re}(z)>0$.
@@
$\mathbb{E}[X^{r}]=\frac{1}{\Gamma(-r)}\int_{0}^{\infty}\frac{1}{t^{r+1}}M_{X}(-t)dt,\qquad(X>0),$